OpenStudy (dls):

The AM and HM of roots of a quadratic equation are 9 and 16/9 then the equation is?

OpenStudy (jiteshmeghwal9):

the real form of a quadratic equation is\[\LARGE{ax^2-(\alpha + \beta)x+c}\]put the roots & solve it

OpenStudy (anonymous):

HM = 2ab/(a+b) AM = a+b/2

OpenStudy (jiteshmeghwal9):

sorry i forgot c= alpha * beta

OpenStudy (anonymous):

1 st find a+b and ab

OpenStudy (anonymous):

AM = a+b/2 = 9 a+b = 18

OpenStudy (jiteshmeghwal9):

\[\LARGE{ax^2-\left( 9+{16 \over 9 }\right)+\left( 9 \times {16 \over 9} \right)}\]

OpenStudy (anonymous):

HM = 2ab/(a+b) = 16/9

OpenStudy (anonymous):

Find ab?

OpenStudy (anonymous):

Then....Form a quad eq..))

OpenStudy (anonymous):

u get ab = 16

OpenStudy (anonymous):

nw fom...quad eq...))

OpenStudy (anonymous):

x^2 - (a+b)x + ab =0

OpenStudy (anonymous):

so. x^2 - 18x + 16=0

OpenStudy (jiteshmeghwal9):

\[\LARGE{ax^2-\left( 81+16 \over 9 \right)+\left( 16 \right)}\]

OpenStudy (jiteshmeghwal9):

gt it @DLS ?? can u d it further ?

OpenStudy (dls):

@Yahoo! thanks a lot ! @jiteshmeghwal9 your way confused me kinda sorry :/ thanks anyway! :)

OpenStudy (jiteshmeghwal9):

np :)

OpenStudy (dls):

thanks to both of you!

OpenStudy (anonymous):

welcome....

OpenStudy (anonymous):

@jiteshmeghwal9 gud..try...... u knw @DLS he is just in 8th std

OpenStudy (anonymous):

nop..@jiteshmeghwal9 .... i said..u r in 8th......@DLS = 11 std

OpenStudy (dls):

@Yahoo! am i in VIII?

OpenStudy (dls):

lol

OpenStudy (jiteshmeghwal9):

ohhh !

OpenStudy (jiteshmeghwal9):

again mistake

OpenStudy (dls):

haha

OpenStudy (anonymous):

hm..hm....i was not knwing that.....))

OpenStudy (jiteshmeghwal9):

hahaha!

OpenStudy (anonymous):

ur question seems to be in higher level..))

OpenStudy (jiteshmeghwal9):

nope i have tried these questions in my N.T.S.E. books :) @Yahoo!

OpenStudy (anonymous):

i think u may be IIT-Aspirant..))) like ur brother @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

no ! my aim is to give CAT exam :)

OpenStudy (anonymous):

hm.. @DLS wat is ur aim..))

OpenStudy (dls):

IIT JEE 2014

OpenStudy (anonymous):

u r in 11

OpenStudy (dls):

yes

OpenStudy (jiteshmeghwal9):

best of luck both of u for ur IIT-JEE :)

OpenStudy (dls):

thanks!

OpenStudy (anonymous):

thxx...

OpenStudy (jiteshmeghwal9):

:)